OpenStudy (anonymous):

Which inequality models this problem? The length of a rectangle is three times its width. If the perimeter is at most 112 centimeters, what is the greatest possible value for the width? A. 2w + 2 • (3w) < 112 B. 2w + 2 • (3w) ≤112 C. 2w + 2 • (3w) > 112 D. 2w + 2 • (3w) ≥112

OpenStudy (texaschic101):

2w + 2(3w) <= 112 ===> answer greatest possible value : 2w + 6w <= 112 8w <= 112 w <= 112/8 w <= 14 ===> greatest possible value for width is 14 cm.

OpenStudy (anonymous):

So is it a b c or d

OpenStudy (texaschic101):

Its B

OpenStudy (anonymous):

ok thanks how do you give meadles?

OpenStudy (texaschic101):

you just did.....all you have to do is click on best response by the person you want to give the medal to :)

OpenStudy (anonymous):

ok